

********************************************************************
UARS WEEKLY STATUS REPORT - 11 AUG 2000
********************************************************************
***********
SOLSTICE
***********
SOLSTICE is in good health, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. The solar data is not affected by the larger than normal
pointing errors since they are within our calibrations. Additionally, we
continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
Level2 processing is halted due to bugs in the OA_SSPP_ATT service. CDHF
and CSC personnel are working on the problem.
***********
SUSIM
***********
Most of the daily data for August 8 were lost due to the UARS OBC ATC
halt. The solar pointing, degraded due to the loss of the star tracker
has improved recently to a few arcmin. Otherwise, SUSIM continues to
gather a normal set of daily solar scans scheduled for periods of
predicted telemetry.
***********
ACRIM
***********
The ACRIM II instrument is performing normally, producing high quality
data. The decreased SSPP pointing accuracy remains well within ACRIM II's
tolerance and no significant degradation of results is anticipated. Data
re-processing is under way using algorithm improvements derived from the
ACRIMSAT/ACRIM III experiment. The long term record of ACRIM II results
will be brought up to date this fall following completion of the re-processing.
***********
PEM
***********
PEM/AXIS-1 was activated on July 27 and deactivated on July 22. The
full PEM/AXIS was activated July 24 and deactivated July 26. The PEM
zenith spectrometers were reactivated in daylight mode on July 21. All of
PEM was off for the July 12 particle event.
At the present time, PEM energy deposition is not being generated.
Recent data configuration changes have affected PEM software causing multiple
failures. These failures are being reviewed by the PEM team. In order to
correct for new telemetry requirements, PEM will need to update processing
code with a patch delivery. The delivery patch will not include changes which
would allow the PEM software to execute on a DEC Alpha machine.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to coordinate with the Flight Operations
Team and to assist the Project Scientist in scheduling of instrument
operations.
The MPG attended a meeting on 10 August 2000 to discuss the progress of
ongoing efforts to improve UARS attitude control using Fixed Head Star
Tracker 1 and the Fine Sun Sensor. Status of planning for future use of a
magnetic field model and the three-axis magnetometer for spacecraft attitude
control was also discussed.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 29 July 2000 (Orbit 48542, GMT Day 211) through 04 August
2000 (Orbit 48646, GMT Day 217).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 212/1325 48565 On
HALOE 214/2349 48601 On
The beta angle ranged between the angles of 71.3 degrees to 48.3 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
+2 degrees on 14 August 2000 (DOY 227).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 28 July 2000 and 04 August 2000 are provided in Appendix D.
The clock error ranged from -13.1 to 12.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report.
A summary is listed below and the details of ALERT and Out-Of-Limits
occurrences are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 02
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 20
¥ WINDII = 01
¥ HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 02
Four (4) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/145 - 2287.497008
· Transponder B: 00/152 - 2287.498920
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/216 show the observatory data loss to be 296 hours, 54
minutes, 26.417 seconds (an increase of 10 minutes 21.712 seconds since the
last report period). This is a 0.39 percent data loss which equals a 99.61
percent data capture for the mission.
The increase of 10 minutes 21.712 seconds occurred over three (3) days:
· 00 minutes 03.056 seconds on Day 210 due to a late acquisition,
· 10 minutes 18.080 seconds on Day 212 due to an incorrect vector,
· 00 minutes 00.576 seconds on Day 213 due to a data hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
UARS continues to operate FHST-1 this week. More updates to the scale factor
were uplinked this week on 31 July, 02 August and 04 August 2000. These
updates are now being provided on Monday, Wednesday and Friday from FDF. A
star intensity calibration was also uplinked on 02 August 2000. The FOT will
continue to perform calibrations so that the FHST-1 is performing at its best
for the yaw maneuver on 14 August 2000. The star tracker is now being cycled
on for one orbit each day to conserve its useful life.
The FOT has delayed the testing of the new TMONs until the FHST-2 anomaly is
resolved. No time frame has been set as to when testing will resume.
The FOT has begun writing the online procedure for safehold should it happen
before the new TMONs are in place or during and OBC halt. This procedure will
be tested on the simulator and put online for immediate execution should an
anomalous condition take place.
MEETINGS
The FOT attended a meeting to discuss FHST-1 on 02 August 2000. The main
action items from this meeting were to begin power cycling the star tracker,
look into how accurate UARS can be with just the TAM and FSS and compare a 5th
order magnetic field model with a 7th order model.
The FOT attended the monthly status meeting on 02 August 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for August 14, 2000 (DO
ght maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 04 Aug 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
48549 211/1146 00-373 OBC Clk Adj fr 20.6 to 41.2 msec/day
48565 212/1325 00-374 WINDII Power ON
48568 212/1825 00-375 OBC Clk Adj fr 41.2 to 20.6 msec/day
48572 213/0126 00-376 OBC Clk Adj fr 20.6 to 0.00 msec/day
48584 213/2011 00-377 FHST1 SCALE FACTOR UPDATE
48598 214/1917 00-379 OBC Clk Adj fr 0.00 to 20.6 msec/day
48601 214/2349 00-380 HALOE Power ON
48605 215/0631 00-378 NBTR NOOPS D-215
48613 215/1832 00-381 NBTR NOOPS D-216
48613 215/1832 00-382 FHST-1 Intensity Calibration
48615 215/2207 00-383 FHST-1 Scale Factor Update
48619 216/0353 00-383 MACS FHST 1 POWER OFF
48620 216/0534 00-381 NBTR NOOPS D-216
48625 216/1414 00-383 MACS FHST-1 POWER ON
48626 216/1555 00-385 MACS FHST-1 POWER OFF
48628 216/1834 00-384 NBTR NOOPS D-217
48626 216/1914 00-386 OBC Clk Adj fr 20.6 to 0.00 msec/day
48634 217/0355 00-387 FHST-1 Power ON
48635 217/0535 00-387 FHST-1 Power OFF
48641 217/1556 00-388 FHST-1 Power ON
48642 217/1736 00-388 FHST-1 Power OFF
48643 217/1837 00-389 NBTR NOOPS D-218
48644 217/2058 00-390 FHST/OBC Scale Factor Update
______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 04 Aug 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
OBSZ1OOL UFLTR Z1 Residual Exceeded 100 Arcsec Continuous
OBSZ2OOL UFLTR Z2 Residual Exceeded 100 Arcsec Continuous
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 04 Aug 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
48549 211/11:46 (29 July 00) from 20.6 to 41.2 msec/day
48568 212/18:34 (30 July 00) from 41.2 to 20.6 msec/day
48572 213/01:33 (30 July 00) from 20.6 to 00.0 msec/day
48598 214/19:17 (01 Aug 00) from 00.0 to 20.6 msec/day
48626 216/19:26 (03 Aug 00) from 20.6 to 00.0 msec/day
_______________________________________________________________________________
UARS BATTERY PERFORMANCE DATA
Week ending 04 Aug 2000
28 July 2000 - GMT Day 210 Beta = 71.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ----------- ------- ------ ------
1 +67.2/+44.8 -2.42/-3.68 +0.0/+0.0 15.5 0.000 0.0
2 +16.8/-11.2 +3.23/+1.66 +17.6/-13.6 27.0 1.028 14.8
3 +22.4/-22.4 +2.60/+1.34 +16.8/-13.6 27.0 1.030 14.5
04 Aug 2000 - GMT Day 217 Beta = 48.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ---------- ------- ------- ------
1 +78.4/+56.0 -3.05/-4.31 +0.0/+0.0 15.5 0.000 0.0
2 +44.8/-33.6 +5.79/+4.19 +27.2/-15.2 25.0 1.020 21.6
3 +22.4/-44.8 +4.51/+3.23 +26.8/-15.2 24.8 1.020 21.2
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 04 Aug 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-034 FOT 48566 212/1547 22739 No Lock/Bad Vector
00-035 FOT 48649 218/0455 22770 Dropouts/Ret. Link
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-033 FOT 48509 208/1958 2723 Late Acq.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-026 FOT see list June 00 14170 Generic Late Acq.
00-032 FHST 48211 188/2242 N/A FHST 2 Anomaly
00-034 FOT 48566 212/1547 22739 No Lock/Bad Vector
00-035 FOT 48649 218/0455 22770 Dropouts/Ret. Link
SEVNTFW00/31